OpenStudy (anonymous):

Simplify. Please be sure to show all of your work.? 92- 5*7 / 2^3 - 7^2-(-6)^2

OpenStudy (anonymous):

Using BEDMAS (Brackets Exponents Division Multiplication Addition Subtraction) as my order of operations. I'm also assuming a linear method so, 2^3 = 8. 2^7 = 128. -6^2 = 36. 5*7 = 35. so 92-35/8-128-36 = 92-4.375-128-36 = -76.375.

OpenStudy (anonymous):

92 - [5*7 / (2^3)] - (7^2) - [(-6)^2] ........ do the exponents first 92 - [5*7 / 8] - (49) - (36) ......... now multiply 92 -[35/8] -49 -36 92 -[4and 3/8] - 49 - 36 .... combine negatives 92 - (89 and 3/8) 2 and 5/8

OpenStudy (anonymous):

is wat i got for some reason